A freight train traveled to Las Vegas and back. On the trip there it traveled 50 mph and on teh return trip it went 40 mph. How long did the trip there take if the return trip took 15 hours?
Answer by Alan3354(30993)

A freight train traveled to Las Vegas and back. On the trip there it traveled 50 mph and on teh return trip it went 40 mph. How long did the trip there take if the return trip took 15 hours?
-----------
On the return, it was 15 hrs @ 40 mi/hr
d = r*t
=========
Find the distance.
For the trip going,
d = r*t
t = d/r